A sample of a gas occupies 5.0 L at STP. What will be the volume if the temperature is set to 320K at a pressure of 1.2 atm? 


1
Expert's answer
2021-10-12T07:48:12-0400

We write mendeleev-Clayperon equation for non-standard conditions


"\frac{P_1V_1}{T_1}=

\frac{P_2V_2}{T_2}." Where "P_1V_1T_1=STP"



"V_2=\frac{1.013×10^3×5×320}{121590×273}=4.883L"